OpenStudy (anonymous):

Solve for x: x - = 8 2 A. x = 17 B. x = 18 C. x = 16 D. x = 12

Nnesha (nnesha):

get rid of the fraction from left side it's x `divided ` by 2 do opposite of mulitplication to cancel out the 2

OpenStudy (yese30):

x=16

OpenStudy (anonymous):

X/2 = 8 X = 8 * 2 X = 16 The option is C

Nnesha (nnesha):

opposite of `division ` typo

OpenStudy (mathstudent55):

@Nnesha You wrote "do opposite of mulitplication to cancel out the 2 " I think you meant: do the opposite of dividing by 2, which is multiply by 2, to cancel out the 2.

OpenStudy (mathstudent55):

I see.

Nnesha (nnesha):

haha

OpenStudy (mathstudent55):